By selling a dinner set for rs 4,320 ,a crockery dealer  made  a profit of 8% .find the cp of dinner set 

Grade:7

1 Answers

Arun
4 years ago
Let cost of dinner set = x
Then profit = (4320- x) *100/x
But in question, profit is 8%
Hence
(4320 -x)*100/x = 8
4320*100 - 100x = 8x
108 x = 432000
Hence x = 432000/108
x = 4000 rs
Hence c.p. is 4000
 
Hope it helps
 
Regards